how much distilled water should be added to 100 ml of a highly concentrated 90% saline solution to reduce the concentration to 40%? what is the volume of the resulting mixture?
Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(21605)

how much distilled water should be added to 100 ml of a highly concentrated 90% saline solution to reduce the concentration to 40%? what is the volume of the resulting mixture?
------------------------
100 ml of 90% salt contains 90 ml of salt.
To add water to make that 90 ml 40% of the solution:
90 ml = 40%
9 ml = 4%
times 25
225 ml = 100%
-------------
125 ml of water has to be added

Answer by RAY100(1637)

Let x = water added
.
x (0%) + 100ml * .9 = (100+x) * .4,,,,,salt balance
.
90 = 40 + .4x
.
-.4x = -50
.
x= 125 ml,,,,h2o added
.
Total volume is 100 + 125 = 225 ml
.
check,,,,100*.9 = 225* .4 = 90,,,,,ok